# GigaHands: A Massive Annotated Dataset of Bimanual Hand Activities

### v1

@@ Abstract
Figure 1: GigaHands is a massive dataset of human bimanual activities with paired text annotations. Each column above shows an activity sequence from the dataset. The dataset covers diverse 3D hand activities, including hand-object interactions (blue) across scales, gestures (orange), and self-interactions (red). Each clip is paired with descriptive text and 51 camera views, enabling radiance field reconstruction. The bottom row show other annotations in the dataset including hand shape, object shape and pose (left half images). The right half images show novel views from dynamic radiance field fitting.
@@ 1 Introduction
We address the data sourcing problem preventing the scaling up of hand datasets by introducing GigaHands, a diverse, massive, and fully-annotated 3D bimanual hand activities dataset (see Figure 1). To our knowledge, GigaHands is the largest bimanual hand activities dataset, with over 183 million unique image frames with two hands each (0.37$\times$109 unique hand poses, hence GigaHands). We used a multi-camera markerless capture system to acquire accurate bimanual hand-object interaction activities, gestures, and self-contacts. We overcome the studio capturing limitation by designing activity elicitation protocols to include as many actions as in the real-world setting. To ensure diversity that reflects the real world, we captured data from 56 subjects interacting with 417 real-world objects. All images in our dataset are fully annotated with: detailed activity text descriptions; 3D hand shape and pose; MANO [romero2022embodied] hand meshes; 3D object shape, pose and appearance; hand/object segmentation masks; 2D/3D hand keypoints; camera pose. We minimized manual annotation effort using an instruct-to-annotate strategy by guiding subjects with detailed instructions to reduce annotation effort post-capture. In total, we collected 34 hours of text-annotated bimanual hand activities, 13k 3D motion sequences, and over 14k post-processed 3D motion clips with 84k detailed atomic-level text description annotations. Our 3D hand motion clips exceed the combined size of all existing 3D bimanual hand activities datasets [liu2022hoi4d, liu2024taco, zhan2024oakink2], and the range of verbs in our text annotations is larger than any other hand dataset, even those captured in the wild [grauman2022ego4d, grauman2024ego].
@@
  • A data acquisition strategy that ensures activity diversity and detailed atomic-level text descriptions while minimizing manual effort. We also built an accurate and fully-automated pipeline for estimating 3D hand/object shape and pose, segmentation, and camera pose.
@@ 2.1 Hand Motion Data Sourcing
Human motion data can be sourced through four primary methods: static poses, synthetic data, in-the-wild settings, and studio environments. Each method presents its own advantages and challenges concerning realism, diversity, accuracy, and practicality.
@@
In-the-wild settings involves using wearable or portable sensors to capture hand interactions from egocentric [pirsiavash2012detecting, damen2018scaling, li2021eye, damen2022rescaling, grauman2022ego4d, liu2022hoi4d], third-person [goyal2017something, miech2019howto100m, shan2020understanding] or both types of views [sigurdsson2018charades, grauman2024ego], yielding realistic and natural motions. However, obtaining reliable 3D motion data for both hands and objects is challenging due to occlusions resulting from limited sensor deployment and accuracy. Additionally, annotating such data is labor-intensive, and certain self-contact motions are difficult to record accurately.
@@ 2.2 Hand Datasets and Annotations
Text annotations. To capture the rich semantic meanings in human hand actions, certain datasets provide text annotations. These annotations may include action types [pirsiavash2012detecting, goyal2017something, dreher2019learning, krebs2021kit, grauman2022ego4d], atomic action descriptions [damen2018scaling, li2021eye, damen2022rescaling, sener2022assembly101, grauman2024ego], activity narration [damen2018scaling, miech2019howto100m, damen2022rescaling, grauman2024ego], activity commentary [grauman2022ego4d, grauman2024ego, zhan2024oakink2, cha2024text2hoi], object affordance [yang2022oakink, jian2023affordpose, zhan2024oakink2, liu2024taco], or body dynamics [zhang2024nl2contact, zhang2024both2hands, yu2024signavatarslargescale3dsign]. Typically, annotations are manually provided, especially when data are collected in uncontrolled, unscripted environments [damen2018scaling, damen2022rescaling, grauman2022ego4d, grauman2024ego]. Even in studio settings, post-processing is often required to extract meaningful action clips [li2021eye, zhan2024oakink2]. We introduce an instruct-and-annotate pipeline that maximizes semantic interpretability and minimizes the annotation effort.
Contact analysis and generation is also important. Contacts are derived from manual annotation [shan2020understanding], video grounding [nagarajan2019grounded], thermal sensors [brahmbhatt2019contactdb, brahmbhatt2020contactpose] or geometry analysis [taheri2020grab, zhu2024contactart, pokhariya2024manus] based on captured data. Our dataset enables us to derive contact regions using geometry analysis on both surface mesh and density fields.
@@ 3 The GigaHands Dataset
Diverse Hand Pose and Motion Variation.
Figure 2 (left) compares hand pose diversity across datasets using t-SNE [van2008visualizing] clustering of 3D keypoints of both hands. Similarly, Figure 2 (middle) shows hand motion diversity through t-SNE clustering of the latent codes of a Variational Autoencoder trained for motion reconstruction. GigaHands exhibits significantly greater diversity than existing datasets in both bimanual hand pose and motion. Additionally, datapoints near the boundary of our dataset seem to correspond to text annotations that are unique in our datasets.
@@ 4 Dataset Acquisition
We present our data acquisition pipeline, which balances realism, diversity, and accuracy while reducing annotation effort. Our pipeline, called Instruct-to-Annotate (see Figure 4), consists of several key components:
  • Instruction elicitation (Section 4.1) for generating instructions to guide participants during data capture.
@@ 4.1 Instruction Elicitation
To ensure diverse hand activities and reduce annotation effort, the Instruct-to-Annotate pipeline starts with an instruction elicitation protocol. We began by sourcing atomic actions from multiple datasets [grauman2022ego4d, grauman2024ego, zhan2024oakink2, liu2024taco], extracting a pool of verbs corresponding to those actions. To elicit subjects to perform these verbs (actions), we manually associated each verb with multiple objects and, with the assistance of an LLM [achiam2023gpt], grouped the verbs and objects into different scenarios such as cooking and eating, office work, crafting, entertainment, and housework. We then structured these scenarios into scenes where the objects could co-occur. Using LLM, we organized the scenes into lists of activities that utilized the objects and verbs in a temporally smooth order and automatically generated detailed instruction scripts. These scripts comprise 5 scenarios, 25 scenes, 191 activities, and 1370 instructions containing a total of 533 verbs. Please see the supp. document for details.
@@ 4.2 Filming
To capture our dataset, we use a custom-designed multi-camera tabletop capture system that consists of 51 RGB cameras uniformly arranged within a cubic capture volume, with each face of the cube containing a 3$\times$3 grid of cameras evenly illuminated by LED lights. Inside the cube, a transparent glass surface serves as a supportive platform for objects. Each camera records at 30fps with a resolution of 1280$\times$720. Cameras are software-synchronized, with the temporal phase misalignment being less than 3 ms. Camera intrinsics and extrinsics are obtained using COLMAP [schoenberger2016sfm, schoenberger2016mvs] aided by fiducial markers.
@@ 4.3 Action Annotation & Augmentation
Though each filmed motion sequence was paired with an instruction, manual annotation was still necessary for two reasons. First, the instructions generated by the LLM occasionally contained inconsistencies or hallucinations. Second, participants might misunderstand the instructions or add extra actions during filming. To encourage subjects to act freely and naturally, we intentionally retained these recordings. Annotators then split these sequences into individual clips and annotated any actions not included in the original instructions. This process resulted in a more accurate dataset with matched (description, clip) pairs. In total, we refined the 13k motion sequences into 14k motion clips.
@@ 5.1 Text-driven Hand Motion Synthesis
We split GigaHands into train, test, and val sets with a ratio of 16:3:1. From among our annotations, we chose $42$ 3D keypoints for both hands as the representation for model training (see the supp. document for an analysis of alternative representations). To evaluate the naturalness, diversity, and alignment of generated motions with textual descriptions, we use metrics from [Guo_2022_CVPR], including R-Precision, Multimodal Distance (MM Dist), Fréchet Inception Distance (FID), Diversity (Div.), and Multimodality (MM.). For computing FID and Div., we train a motion autoencoder to define a compressed motion space. For R-Precision, MM Dist, and MM., we employ contrastive learning to create a joint motion-text embedding space, following [Guo_2022_CVPR]. Feature extractors are trained independently for each dataset and subset.
@@
upper bound 64.4 86.2 94.2 2.86 0.045 14.2 -
TACO [liu2024taco] 19.6 37.3 52.3 9.41 9.5 11.1 6.01
@@
OakInk2 [zhan2024oakink2] 20.3 39.1 55.1 5.68 12.6 4.17 2.45
@@ Results.
We report text to hand motion synthesis performance using the TM2T [chuan2022tm2t] backbone in Table 2 (see the supp. document for more results on transformer [zhang2023generating] and diffusion [tevet2023human] architectures). Since the evaluation embedding spaces are trained on different test sets, we also evaluate the metrics on ground truth test data to indicate upper-bound performance for each test set. Models trained on GigaHands outperform others on all metrics except MM Dist., with higher R-Precision due to better text-motion alignment. Compared to OakInk2 (which also includes textual descriptions for hand motions), GigaHands achieves significantly better FID, Diversity, and Multimodality scores, likely due to its greater dataset diversity. Figure 5 shows hand motions generated by TM2T trained with GigaHands using text inputs from the test sets of GigaHands, OakInk and TACO. Even without object geometry input, our model generates reasonable hand shapes and poses for object manipulation, benefiting from the diversity of objects in our dataset. Our model can also generate reasonable motion using text from other datsasets, showing the comprehensiveness of GigaHands.
@@
Figure 6 illustrates the impact of dataset size on motion reconstruction and text-to-motion generation. Based on the TM2T architecture, we train a motion VQ-VAE for reconstruction and an attentive GRU-based text-to-motion model for generation using 10%, 20%, 50%, 80%, and 100% of the training set, evaluating on the same test set. We evaluate FID, MM Dist., and Top 1 & Top 3 accuracies. Most metrics continually improve with larger datasets, demonstrating the value of larger-scale data.
@@ 5.2 Hand Motion Captioning
In-the-Wild 3D Hand Motion Captioning
@@ 6 Conclusion
We present GigaHands, a massive annotated dataset of bimanual hand activities. The dataset contains 14k motion clips with 3D hand and object motions from 56 subjects interacting with 417 real-world objects, all captured from 51 camera views. It provides 183 million unique image frames and 84k textual descriptions, enabling a wide range of applications. We demonstrated how the scale and diversity of the dataset benefit text-driven motion synthesis and motion captioning, both within the dataset and on in-the-wild data. Additionally, we have shown that GigaHands enables dynamic radiance field reconstruction, opening possibilities for downstream tasks.
Limitations and Future Directions. Despite its scale and diversity, GigaHands has limitations. The studio setting confines data collection to a limited space, making it challenging to accurately capture motions that require larger environments. While we can track rigid objects and object parts, fully automatic tracking of articulated and non-rigid objects remains challenging. Advancements in markerless tracking techniques and more flexible representations are needed to address this issue. Moreover, although we have showcased applications in motion synthesis and captioning, further research could explore how the dataset’s scale and diversity can enhance robotic manipulation and human-computer interaction tasks. Expanding the dataset’s applicability to these areas could significantly advance the development of more natural and capable AI systems.

### v3

@@ Abstract
Figure 1: GigaHands is a massive dataset of human bimanual activities with paired text annotations. Each column above shows an activity sequence from the dataset. The dataset covers diverse 3D hand activities, including hand-object interactions (blue) across object scales, gestures (orange), and self-interactions (red). Each clip is paired with descriptive text and 51 camera views, enabling radiance field reconstruction. The bottom row show other annotations in the dataset including hand shape, object shape and pose (left half images). The right half images show novel views from dynamic radiance field fitting.
@@ 1 Introduction
[…]
@@
  • A procedural data acquisition strategy that ensures activity diversity and detailed atomic-level text descriptions while minimizing manual effort. We also built an accurate and fully-automated pipeline for estimating 3D hand/object shape and pose, segmentation, and camera pose.
@@ 2.1 Hand Motion Data Sourcing
Hand motion data has four primary sources, and each presents its own advantages and challenges concerning realism, diversity, accuracy, and practicality.
@@
Real-world in-the-wild settings involves using wearable or portable sensors to capture hand interactions from egocentric [pirsiavash2012detecting, damen2018scaling, li2021eye, damen2022rescaling, grauman2022ego4d, liu2022hoi4d], third-person [goyal2017something, miech2019howto100m, shan2020understanding] or both types of views [sigurdsson2018charades, grauman2024ego], yielding realistic and natural motions. However, obtaining reliable 3D motion data for both hands and objects is challenging due to occlusions resulting from limited sensor deployment and accuracy. Additionally, annotating such data is labor-intensive, and certain self-contact motions are difficult to record accurately.
@@ 2.2 Hand Datasets and Annotations
Text annotations. To capture the rich semantic meanings in human hand actions, certain datasets provide text annotations. These annotations may include action types [pirsiavash2012detecting, goyal2017something, dreher2019learning, krebs2021kit, grauman2022ego4d], atomic action descriptions [damen2018scaling, li2021eye, damen2022rescaling, sener2022assembly101, grauman2024ego], activity narration [damen2018scaling, miech2019howto100m, damen2022rescaling, grauman2024ego], activity commentary [grauman2022ego4d, grauman2024ego, zhan2024oakink2, cha2024text2hoi], object affordance [yang2022oakink, jian2023affordpose, zhan2024oakink2, liu2024taco], or body dynamics [zhang2024nl2contact, zhang2024both2hands, yu2024signavatarslargescale3dsign]. Typically, annotations are manually provided, especially when data are collected in uncontrolled, unscripted environments [damen2018scaling, damen2022rescaling, grauman2022ego4d, grauman2024ego]. Even in studio settings, post-processing is often required to extract meaningful action clips [li2021eye, zhan2024oakink2]. We introduce a procedural instruct-and-annotate pipeline that maximizes semantic interpretability and minimizes the annotation effort.
Contact are derived from manual annotation [shan2020understanding], video grounding [nagarajan2019grounded], thermal sensors [brahmbhatt2019contactdb, brahmbhatt2020contactpose] or geometry analysis [taheri2020grab, zhu2024contactart, pokhariya2024manus] based on captured data. Our dataset enables us to derive contact regions using geometry analysis on both surface mesh and density fields.
@@ 3 The GigaHands Dataset
Diverse Hand Pose and Motion.
Figure 2 (left) compares hand pose diversity across datasets using t-SNE [van2008visualizing] clustering of 3D keypoints of both hands. Similarly, Figure 2 (middle) shows hand motion diversity through t-SNE clustering of the latent codes of a Variational Autoencoder trained for motion reconstruction. GigaHands exhibits significantly greater diversity than existing datasets in both bimanual hand pose and motion. Additionally, datapoints near the boundary of our dataset seem to correspond to text annotations that are unique in our datasets, demonstrating that diverse verbs contribute to diverse poses and motions.
@@ 4 Dataset Acquisition
We present our data acquisition pipeline, which balances realism, diversity, and accuracy while reducing annotation effort. Our procedural pipeline, called Instruct-to-Annotate (see Figure 4), consists of several key components:
  • Procedural instruction elicitation (Section 4.1) for generating instructions to guide participants during data capture.
@@ 4.1 Instruction Elicitation
To ensure diverse hand activities and reduce annotation effort, the Instruct-to-Annotate pipeline starts with a procedural instruction elicitation protocol. We began by sourcing atomic actions from multiple datasets [grauman2022ego4d, grauman2024ego, zhan2024oakink2, liu2024taco], extracting a pool of verbs corresponding to those actions. To elicit subjects to perform these verbs (actions), we manually associated each verb with multiple objects and, with the assistance of an LLM [achiam2023gpt], grouped the verbs and objects into different scenarios such as cooking and eating, office work, crafting, entertainment, and housework. We then structured these scenarios into scenes where the objects could co-occur. Using LLM, we organized the scenes into lists of activities that utilized the objects and verbs in a temporally smooth order and automatically generated detailed instruction scripts. These scripts comprise 5 scenarios, 25 scenes, 191 activities, and 1370 instructions containing a total of 533 verbs. Please see the supp. document for details.
@@ 4.2 Filming
[…]
@@ 4.3 Action Annotation & Augmentation
Though each filmed motion sequence was paired with an instruction, manual annotation was still necessary for two reasons. First, the instructions generated by the LLM occasionally contained inconsistencies or hallucinations. Second, participants might misunderstand the instructions or add unintentional actions during filming. To encourage subjects to act freely and naturally, we intentionally retained these recordings. Annotators then split these sequences into individual clips and annotated any actions not included in the original instructions. This process resulted in a more accurate dataset with matched (description, clip) pairs. In total, we refined the 13k motion sequences into 14k motion clips.
@@ 5.1 Text-driven Hand Motion Synthesis
We split GigaHands into train, test, and val sets with a ratio of 16:3:1. From among our annotations, we chose $4242$ 3D keypoints for both hands as the representation for model training (see the supp. document for an analysis of alternative representations). To evaluate the naturalness, diversity, and alignment of generated motions with textual descriptions, we use metrics from [Guo_2022_CVPR], including R-Precision, Multimodal Distance (MM Dist), Fréchet Inception Distance (FID), Diversity (Div.), and Multimodality (MM.). For computing FID and Div., we train a motion autoencoder to define a compressed motion space. For R-Precision, MM Dist, and MM., we employ contrastive learning to create a joint motion-text embedding space, following [Guo_2022_CVPR]. Feature extractors are trained independently for each dataset and subset.
@@
upper bound 64.4 86.2 89.4 2.86 0.045 14.2 -
TACO [liu2024taco] 18.9 37.7 52.9 7.39 11.0 11.1 6.83
@@
OakInk2 [zhan2024oakink2] 17.9 31.7 47.9 7.75 19.6 6.88 3.45
@@ Results.
We report text to hand motion synthesis performance using the T2M-GPT [zhang2023generating] backbone in Table 2 (see the supp. document for more results on GRU-based [chuan2022tm2t] and diffusion [tevet2023human] architectures). Since the evaluation embedding spaces are trained on different test sets, we also evaluate the metrics on ground truth test data to indicate upper-bound performance for each test set. Models trained on GigaHands outperform others on all metrics except MM Dist., with higher R-Precision due to better text-motion alignment. Compared to OakInk2 (which also includes textual descriptions for hand motions), GigaHands achieves significantly better FID, Diversity, and Multimodality scores, likely due to its greater dataset diversity. Figure 5 shows hand motions generated by T2M-GPT trained with GigaHands using text inputs from the test sets of GigaHands, OakInk and TACO. Even without object geometry input, our model generates reasonable hand shapes and poses for object manipulation, benefiting from the diversity of objects in our dataset. Our model can also generate reasonable motion using text from other datsasets, showing the comprehensiveness of GigaHands.
@@
Figure 6 illustrates the impact of dataset size on motion reconstruction and text-to-motion generation. Based on the T2M-GPT architecture, we train a motion VQ-VAE for reconstruction and a generative pretrained transformer model for text-to-motion generation using 10%, 20%, 50%, 80%, and 100% of the training set, evaluating on the same test set. We evaluate FID, MM Dist., and Top 1 & Top 3 accuracies. Most metrics continually improve with larger datasets, demonstrating the value of larger-scale data.
@@ 5.2 Hand Motion Captioning
3D Hand Motion Captioning for Unlabeled Datasets
@@ 6 Conclusion
We present GigaHands, a massive annotated dataset of bimanual hand activities. The dataset contains 14k motion clips with 3D hand and object motions from 56 subjects interacting with 417 real-world objects, all captured from 51 camera views. It provides 183 million unique image frames and 84k textual descriptions, enabling a wide range of applications. We demonstrated how the scale and diversity of the dataset benefit text-driven motion synthesis and motion captioning, both within the dataset and on other data. Additionally, we have shown that GigaHands enables dynamic radiance field reconstruction, opening possibilities for downstream tasks.
Limitations and Future Directions. Despite its scale and diversity, GigaHands has limitations. The studio setting confines data collection to a limited space, making it challenging to accurately capture motions that require larger environments. While we can track rigid objects and object parts, fully automatic tracking of articulated and non-rigid objects remains challenging. Moreover, although we have showcased applications in motion synthesis and captioning, further research could explore how the dataset’s scale and diversity can enhance robotic manipulation and human-computer interaction tasks.